# Elementary Analysis of Policy Gradient Methods

### v1

@@ 1 Introduction
Reinforcement Learning (RL) has achieved great success in the fields of both science and engineering [6, 8, 19, 25]. As the fundamental model for RL, Markov Decision Process (MDP) can be represented as a tuple ${\cal M}({\cal S},{\cal A},P,r,\gamma)$, where ${\cal S}$ is the state space and ${\cal A}$ is the action space, $P(s^{\prime}|s,a)$ is the transition probability from state $s$ to state $s^{\prime}$ under action $a$, $r:{\cal S}\times{\cal A}\rightarrow\mathbb{R}$ is the reward function, and $\gamma\in[0,1)$ is the discounted factor. In this paper, we focus on the finite MDP setting, i.e., $|{\cal S}|<\infty$ and $|{\cal A}|<\infty$, and also assume $r\in[0,1]$ for simplicity. Letting
@@
The overall goal of RL is to seek an optimal policy to maximize the state value, which can be expressed as
@@
Policy optimization refers to a family of effective methods for solving challenging RL problems. In contrast to value-based methods such as value iteration and policy iteration, policy optimization methods conduct direct search in the policy space through the parameterization of policies. More precisely, given a parameterized policy class $\{\pi_{\theta}:\mathbb{R}^{d}\to\Pi\}$, the state value in (1.1) becomes a function of the parameters, denoted $V^{\pi_{\theta}}(\mu)$. Thus, the problem of seeking an optimal policy over the parameterized policy class turns to be a finite dimensional problem, and different kinds of optimization can be applied to solve this problem. Maybe the simplest one among them is the policy gradient method [28, 26],
@@
where $\nabla_{\theta}V^{\pi_{\theta}}\left(\mu\right)$ is the policy gradient of $V^{\pi_{\theta}}(\mu)$. Different parameterization methods lead to different policy gradient methods, for instance projected policy gradient (PPG) under simplex or direct parameterization and softmax policy gradient (PG) under the softmax parameterization. More details will be given in later sections. Natural policy gradient (NPG, [9]) is an important variant of policy gradient methods, which searches along a preconditioned policy gradient direction based on the Fisher information matrix. Other important variants of policy optimization methods include the trust region policy optimization (TRPO) method [22] and the proximal policy optimization (PPO) method [23], which are closely related to NPG.
@@
where $\tau>0$ is the regularization parameter and $\mathcal{H}(p)=-\sum_{a}p_{a}\log p_{a}$ is the entropy of a probability distribution. Intuitively, larger entropy represents more possibilities. The policy gradient method for
@@
is overall similar to that in (1.2), but with $\nabla_{\theta}V^{\pi_{\theta}}\left(\mu\right)$ being replaced by $\nabla_{\theta}V^{\pi_{\theta}}_{\tau}\left(\mu\right)$. Moreover, NPG can be similarly developed for the entropy regularization scenario.
@@ 1.1 Related Work
The convergence analyses of policy optimization methods have received a lot of attention recently under various settings, for example from the tabular MDP setting to the setting that incorporates the functions approximation, from the setting where exact gradient information is assumed to the setting that involves sample-based estimators. An exhaustive review about this is beyond the scope of this paper. Thus, we mainly focus on the convergence results of different exact policy gradients in the tabular setting, which are mostly related to our work.
@@
For exact policy gradient methods under softmax parameterization, the global convergence of softmax PG is established in [1] provided the step size $\eta$ satisfies $\eta\leq 1/L$, where $L$ is the smoothness constant of the value function. By utilizing the gradient dominant property of the value function, the $O(1/k)$ sublinear convergence of softmax PG is established in [18]. In addition, it is shown in [16] that softmax PG can also achieve linear convergence using the adaptive step sizes based on the geometric information. In the presence of regularization, it is shown in [1] that softmax PG with log barrier regularization attain an $O(1/\sqrt{k})$ convergence rate, while the linear convergence is established for softmax PG with entropy (simply referred to as entropy softmax PG later) in [18].
The $O(1/\sqrt{k})$ convergence rate of NPG under softmax parameterization is obtained in [24], and this rate has been improved to $O(1/k)$ in [1]. The fast local linear convergence rate of softmax NPG is established in [11] based on the contraction of the non-optimal probability measure at each state. For softmax NPG with entropy regularization (simply referred to as entropy softmax NPG later), the linear convergence is established in [5] by expressing the update into a form for which the $\gamma$-contraction property of the Bellman operator can be used. This result has been extended to more general regularizers in [31]. Note that softmax NPG can be cast as a special policy mirror descent (PMD) method with KL divergence, and the convergence of PMD with and without regularizers has also been studied. The $O(1/k)$ sublinear convergence of PMD is established in [29] for any constant step size which generalizes the analysis in [1] for softmax NPG based on the three-point lemma. In [29], the linear convergence is also provided for the non-adaptive geometrically increasing step sizes, and the $\gamma$-rate convergence of PMD under adaptive step size is established in [7]. The analyses in [29] have extended to log-linear policy and the more general function approximation scenario in [30, 2]. The convergence of PMD with general convex and strongly convex regularizers is studied in [12] and it is shown that PMD with strongly convex regularizer can achieve a linear convergence. By adopting a homotopic technique using diminishing regularization combined with increasing step sizes, the linear convergence of PMD with entropy regularization is established in [14].
@@ 1.2 Main Contributions
Despite the recent intensive investigations, there are still a few important problems to be addressed on the convergence of the basic policy optimization methods, including projected policy gradient (PPG), softmax PG and NPG in the both the non-entropy and entropy cases. A thorough treatment will be given towards this line of research in this paper. The main contributions are summarized as follows:
@@
  • New local quadratic convergence of soft PI and tight local linear convergence of entropy softmax NPG. Soft PI can be viewed as the limit of entropy softmax NPG as the step size $\eta\to\infty$. A new local quadratic convergence rate in the overall form $\gamma^{\displaystyle 2^{k-k_{0}}}$ has been established, see Theorem 6.13. Compared with the result in [5], our result only essentially relies on the discount factor $\gamma$ and does not require an additional assumption on the stationary distribution under the optimal policy. As mentioned earlier, global linear convergence of entropy softmax NPG has also been established in [5]. However, as $\eta\rightarrow\infty$, the limit of the rate is a fixed constant, which cannot reflect the fact that entropy softmax NPG should converge faster (locally) as $\eta$ becomes larger since the algorithm tends to be soft PI and the latter one enjoys a local quadratic convergence rate. Motivated by this observation, a local linear convergence rate in the form of $\left(\frac{1}{(\eta\tau+1)^{2}}\right)^{k-k_{0}}$ is established. It is meanwhile shown that this rate is indeed tight.
@@ 1.3 Outline of This Paper
The rest of this paper is organized as follows. In Section 2 we present a general analysis framework for the policy gradient methods in the non-entropy case, which will used in the subsequent sections. Some of the results can be easily extended to handle the entropy case, and those needed will be mentioned later in Section 6. The new convergence results for PPG, softmax PG and softmax NPG are presented in Sections 3, 4, and 5, respectively. The discussion of softmax PG and softmax NPG in the entropy regularized case is provided in Section 6. This paper is concluded with a few future directions in Section 7.
@@ 2.1 Preliminaries
For an arbitrary vector $V\in\mathbb{R}^{|{\cal S}|}$, the Bellman operator induced by policy $\pi$ is defined as
@@
Let $V^{*}$ and $Q^{*}$ be the optimal state and state-action values, with the corresponding optimal policy denoted by $\pi^{*}$ (may be non-unique), and define $A^{*}(s,a)=Q^{*}(s,a)-V^{*}(s)$. It is well-known that $V^{*}$ satisfies the Bellman optimality equation: ${\cal T}V^{*}=V^{*}$. It is also noted that both ${\cal T}^{\pi}$ and ${\cal T}$ are $\gamma$-contraction under the infinity norm:
@@
Given a state value $V^{k}$ (short for $V^{\pi^{k}}$) associated with policy $\pi^{k}$, policy iteration (PI) attempts to obtain a new policy $\pi^{k+1}$ through greedy improvement:
@@
Before presenting the next lemma, further define ${\cal A}^{*}_{s}$ as the optimal action set of state $s$,
@@
where $d_{\rho}^{\pi^{k,*}}$ is the visitation measure corresponding to certain optimal polity $\pi^{k,*}$, and $\mathcal{T}^{*}$ is the Bellman operator associated with $\pi^{k,*}$. Unless stated otherwise, we will assume the same optimal policy is used across all the $k$ most of the times, and in this case $d_{\rho}^{\pi^{k,*}}$ is simplified to $d_{\rho}^{*}$. It follows from the performance difference lemma that
@@
and thus ${\cal L}_{k}^{*}$ is irrelevant to the choice of the optimal policy. In addition the following bounds hold for $\mathcal{L}^{*}_{k}$, whose proof can be found in Appendix A.
@@ 2.3 Global Linear Convergence Analysis
which completes the proof.
@@
$\forall\,s:\quad\sum_{a}\pi^{k+1}_{s,a}A^{k}_{s,a}\geq C\,\max_{a}A^{k}_{s,a}-\varepsilon_{k}\quad\mbox{and}\quad\sum_{t=0}^{k-1}\frac{\varepsilon_{t}}{(1-(1-\gamma)C)^{t+1}}\leq c_{0}.$ (2.7)
@@
gives the $\gamma$-rate convergence of PMD which is established in [7].
@@ 4 Softmax Policy Gradient
where $Z_{s}^{k}=\sum_{a^{\prime}}\pi_{s,a^{\prime}}^{k}\exp\left(\eta_{s}^{k}\,\pi^{k}_{s,a^{\prime}}A^{k}_{s,a^{\prime}}\right)$ is the normalization factor. For ease of notation, we will denote $\pi^{k}_{s,a}A^{k}_{s,a}$ by $\hat{A}^{k}_{s,a}$ in this section. It is evident that
@@
The $O(1/k)$ sublinear convergence has been established for softmax PG with constant step size in [18]. The analysis therein is essentially based on the smoothness and the gradient domination property of the value function, and thus requires the step size to be sufficiently small. In addition, an adaptive step size based on the norm of the gradient has been proposed in [16], with which softmax PG can achieve linear convergence. In this section, we will extend the $O(1/k)$ sublinear convergence of softmax PG to any constant step size, and a new adaptive step size selection rule will also be introduced.
In contrast to the optimization analysis framework adopted in [18], the idea of analysis here is more elementary and similar to the one used in Section 3. That being said, the extension is by no means trivial since the update of softmax PG is substantially different with that of PPG. In particular, we will establish the improvement lower bound of $\mathcal{T}^{k+1}V^{k}(s)-V^{k}(s)$ for softmax PG in terms of a function of $\max_{a}|\hat{A}^{k}_{s,a}|$ in contrast to $\max_{a}A^{k}_{s,a}$ due to the existence of $\pi_{s,a}^{k}$ in the exponent in (4.2). Moreover, as noted after Lemma 3.1, the establishment of the improvement lower bound in terms of $\max_{a}A^{k}_{s,a}$ for PPG relies heavily on the explicit formula for the projection which is not applicable for softmax PG. Instead, a complete different route is adopted to establish the improvement lower bound for softmax PG. Moreover, the improvement upper bound is also established which enables us to show that the $O(1/k)$ sublinear convergence rate is tight for softmax PG for any constant step size.
@@ 4.1 Improvement Lower and Upper Bounds
We begin with a novel identity that plays a key in establishing the lower and upper improvement bounds.
@@
Noticing that $\max_{a}|\hat{A}^{k}_{s,a}|\leq\frac{1}{1-\gamma}$, one further has
$\sum_{a}\pi^{k+1}_{s,a}A^{k}_{s,a}\leq\left(\exp\left(\frac{2\,\eta_{k}}{(1-\gamma)^{2}}\right)-1\right)(1-\gamma)|\mathcal{A}|\left(\max_{a}|\hat{A}^{k}_{s,a}|\right)^{2},$
since
$\exp\left(\frac{2\eta_{k}}{1-\gamma}x\right)-1\leq\left(\exp\left(\frac{2\eta_{k}}{(1-\gamma)^{2}}\right)-1\right)(1-\gamma)\,x,\quad\mbox{for }x\in\left[0,\frac{1}{1-\gamma}\right].$
@@ 4.2 Convergence Results
We first present the global convergence of softmax PG for any step size sequence $\eta_{k}$ as long as $\inf_{k}\eta_{k}\geq\alpha$ for some $\alpha>0$. The proof overall follows the one for [1] and is presented in Appendix B where we only point out how to generalize the proof in [1] to any step size sequence.
@@
In particular, if the optimal policy is unique, i.e., $|\mathcal{A}_{s}^{*}|=1$, then
@@
$V^{*}(\rho)-V^{k}(\rho)\leq\frac{1}{k}\frac{1}{(1-\gamma)^{3}}\frac{2|\mathcal{A}|}{\kappa^{2}}\left\|\frac{d_{\rho}^{*}}{\rho}\right\|_{\infty}\left\|\frac{1}{\mu}\right\|_{\infty}.$
@@
The proof is completed by using Theorem 2.11 noting that $\left\|\frac{d^{\xi^{k}}_{\rho}}{\rho}\right\|^{-1}\geq\tilde{\rho}$. ∎
@@
It follows easily from (4.5) and the performance difference lemma that $\sum_{a\in\mathcal{A}}{\pi_{s,a}^{k+1}A_{s,a}^{k}}\geq 0$ and $\underset{k\rightarrow+\infty}{\lim}|\hat{A}_{s,a}^{k}|=0$. Thus, $\eta_{k}$ is lower bounded. The application of Theorem 4.4 also implies the global convergence of softmax PG with adaptive step size and $\kappa>0$. It is worth noting that the adaptive step size $\eta_{k}={\eta}/{\left\|\nabla_{\theta}V^{\pi_{\theta_{k}}}\left(\mu\right)\right\|_{2}}$ is considered in [17] and the following linear convergence rate has been established for softmax PG,

### v2

@@ Abstract
Keywords. Reinforcement learning, policy gradient methods, entropy regularization, sublinear convergence, linear convergence, quadratic convergence
@@ 1 Introduction
Reinforcement Learning (RL) has achieved great success in the fields of both science and engineering, see for example [6, 8, 19, 25]. As the fundamental model for RL, Markov Decision Process (MDP) can be represented as a tuple ${\cal M}({\cal S},{\cal A},P,r,\gamma)$, where ${\cal S}$ is the state space and ${\cal A}$ is the action space, $P(s^{\prime}|s,a)$ is the transition probability from state $s$ to state $s^{\prime}$ under action $a$, $r:{\cal S}\times{\cal A}\rightarrow\mathbb{R}$ is the reward function, and $\gamma\in[0,1)$ is the discounted factor. In this paper, we focus on the finite MDP setting, i.e., $|{\cal S}|<\infty$ and $|{\cal A}|<\infty$, and also assume $r\in[0,1]$ for simplicity. Letting
@@
The overall goal of RL is to seek an optimal policy to maximize the state values, which can be expressed as
@@
Policy optimization refers to a family of effective methods for solving challenging RL problems. In contrast to value-based methods such as value iteration and policy iteration, policy optimization methods conduct direct search in the policy space through the parameterization of policies. More precisely, given a parameterized policy class $\{\pi_{\theta}:\mathbb{R}^{d}\to\Pi\}$, the state value in (1.1) becomes a function of the parameters, denoted $V^{\pi_{\theta}}(\mu)$. Thus, the problem of seeking an optimal policy over the parameterized policy class turns to be a finite dimensional optimization problem, and different kinds of optimization methods can be applied to solve this problem. Maybe the simplest one among them is the policy gradient method [28, 26],
@@
where $\nabla_{\theta}V^{\pi_{\theta}}\left(\mu\right)$ is the policy gradient of $V^{\pi_{\theta}}(\mu)$. Different parameterization methods lead to different policy gradient methods, for instance projected policy gradient (PPG) under the simplex or direct parameterization and softmax policy gradient (PG) under the softmax parameterization. More details will be given in later sections. Natural policy gradient (NPG, [9]) is an important variant of policy gradient methods, which searches along a preconditioned policy gradient direction based on the Fisher information matrix. Other important variants of policy optimization methods include trust region policy optimization (TRPO) [22] and proximal policy optimization (PPO) [23], which are closely related to NPG.
@@
where $\tau>0$ is the regularization parameter and $\mathcal{H}(p)=-\sum_{a}p_{a}\log p_{a}$ is the entropy of a probability distribution. Intuitively, larger entropy represents more possibilities. The policy gradient method for the entropy regularized policy optimization problem
@@
is overall similar to that in (1.2), but with $\nabla_{\theta}V^{\pi_{\theta}}\left(\mu\right)$ being replaced by $\nabla_{\theta}V^{\pi_{\theta}}_{\tau}\left(\mu\right)$. Moreover, NPG can be similarly developed for the entropy regularized scenario.
@@ 1.1 Related Work
The convergence analyses of policy optimization methods have received a lot of attention recently under various settings, for example from the tabular MDP setting to the setting that incorporates function approximations, from the setting where exact gradient information is assumed to the setting that involves sample-based estimators. An exhaustive review about this is beyond the scope of this paper. Thus, we mainly focus on the convergence results of different exact policy gradients in the tabular setting, which are mostly related to our work.
@@
For exact policy gradient methods under softmax parameterization, the global convergence of softmax PG is established in [1] provided the step size $\eta$ satisfies $\eta\leq 1/L$, where $L$ is the smoothness constant of the value function. By utilizing the gradient dominant property of the value function, the $O(1/k)$ sublinear convergence of softmax PG is established in [18]. In addition, it is shown in [16] that softmax PG can also achieve linear convergence using the adaptive step sizes based on the geometric information. In the presence of regularization, it is shown in [1] that softmax PG with log barrier regularization attain an $O(1/\sqrt{k})$ convergence rate, while the linear convergence is established for softmax PG under entropy regularization (simply referred to as entropy softmax PG later) in [18].
The $O(1/\sqrt{k})$ convergence rate of NPG under softmax parameterization is obtained in [24], and this rate has been improved to $O(1/k)$ in [1]. The fast local linear convergence rate of softmax NPG is established in [11] based on the contraction of the non-optimal probability measure at each state. For softmax NPG with entropy regularization (simply referred to as entropy softmax NPG later), the linear convergence is established in [5] by expressing the update into a form for which the $\gamma$-contraction property of the Bellman operator can be used. This result has been extended to more general regularizers in [31]. Note that softmax NPG can be cast as a special policy mirror descent (PMD) method with KL divergence, and the convergence of PMD with and without regularizers has also been studied. The $O(1/k)$ sublinear convergence of PMD is established in [29] for any constant step size which generalizes the analysis in [1] for softmax NPG based on the three-point lemma. In [29], the linear convergence is also provided for the non-adaptive geometrically increasing step sizes. The $\gamma$-rate convergence of PMD under the adaptive step sizes is established in [7]. The analyses in [29] have been extended to the log-linear policy and the more general function approximation scenario in [30, 2]. The convergence of PMD with general convex and strongly convex regularizers is studied in [12] and it is shown that PMD with strongly convex regularizer can achieve a linear convergence. By adopting a homotopic technique using diminishing regularization combined with increasing step sizes, the linear convergence of PMD with entropy regularization is established in [14].
@@ 1.2 Main Contributions
Despite the recent intensive investigations, there are still a few important issues to be addressed on the convergence of the very basic policy optimization methods, including projected policy gradient (PPG), softmax PG and NPG in the both the non-entropy and entropy cases. A thorough treatment will be given towards this line of research in this paper. The main contributions are summarized as follows:
@@
  • New local quadratic convergence rate of soft PI and tight local linear convergence rate of entropy softmax NPG. Soft PI can be viewed as the limit of entropy softmax NPG as the step size approaches infinity. A new local quadratic convergence rate in the form of $\gamma^{\displaystyle 2^{k-k_{0}}}$ has been established, see Theorem 6.13. Compared with the result in [5], our result only essentially relies on the discount factor $\gamma$ and does not require an additional assumption on the stationary distribution under the optimal policy. As mentioned earlier, the global linear convergence of entropy softmax NPG has also been established in [5]. However, as $\eta\rightarrow\infty$, the limit of the rate is a fixed constant, which cannot reflect the fact that entropy softmax NPG should converge faster (locally) as $\eta$ becomes larger since the algorithm tends to be soft PI and the latter one enjoys a local quadratic convergence rate. Motivated by this observation, a local linear convergence rate in the form of $\left(\frac{1}{(\eta\tau+1)^{2}}\right)^{k-k_{0}}$ is established. It is meanwhile shown that this rate is indeed tight.
@@ 1.3 Outline of This Paper
The rest of this paper is organized as follows. In Section 2 we present a general analysis framework for the policy gradient methods in the non-entropy case, which will be used in the subsequent sections. Some of the results can be easily extended to handle the entropy case, and those needed will be mentioned later in Section 6. The new convergence results for PPG, softmax PG and softmax NPG are presented in Sections 3, 4, and 5, respectively. The discussion of softmax PG and softmax NPG in the entropy regularized case is provided in Section 6. This paper is concluded with a few future directions in Section 7.
@@ 2.1 Preliminaries
For an arbitrary vector $V\in\mathbb{R}^{|{\cal S}|}$, the Bellman operator induced by a policy $\pi$ is defined as
@@
Let $V^{*}$ and $Q^{*}$ be the optimal state and action values, with the corresponding optimal policy denoted by $\pi^{*}$ (may be non-unique), and define $A^{*}(s,a)=Q^{*}(s,a)-V^{*}(s)$. It is well-known that $V^{*}$ satisfies the Bellman optimality equation: ${\cal T}V^{*}=V^{*}$. It is also noted that both ${\cal T}^{\pi}$ and ${\cal T}$ are $\gamma$-contraction under the infinity norm:
@@
Given a state value $V^{k}$ (short for $V^{\pi^{k}}$) associated with policy $\pi^{k}$, policy iteration (PI) attempts to obtain a new policy $\pi^{k+1}$ through the greedy improvement:
@@
Before presenting the next lemma, further define ${\cal A}^{*}_{s}$ as the optimal action set at state $s$,
@@
where $d_{\rho}^{\pi^{k,*}}$ is the visitation measure corresponding to certain optimal polity $\pi^{k,*}$ chosen at the $k$-th iteration, and $\mathcal{T}^{*}$ is the Bellman operator associated with $\pi^{k,*}$. Unless stated otherwise, we will assume the same optimal policy is used across all the $k$ most of the times, and in this case $d_{\rho}^{\pi^{k,*}}$ is simplified to $d_{\rho}^{*}$. It follows from the performance difference lemma that
@@
and thus ${\cal L}_{k}^{*}$ is irrelevant to the choice of the optimal policy. In addition, the following bounds hold for $\mathcal{L}^{*}_{k}$, whose proof can be found in Appendix A.
@@ 2.3 Global Linear Convergence Analysis
which completes the proof. ∎
@@
[…]
@@
gives the $\gamma$-rate convergence of PMD established in [7].
@@ 4 Softmax Policy Gradient
where $Z_{s}^{k}=\sum_{a^{\prime}}\pi_{s,a^{\prime}}^{k}\exp\left(\eta_{s}^{k}\,\pi^{k}_{s,a^{\prime}}A^{k}_{s,a^{\prime}}\right)$ is the normalization factor. For ease of notation, we will let $\hat{A}^{k}_{s,a}=\pi^{k}_{s,a}A^{k}_{s,a}$ in this section. It is evident that
@@
The $O(1/k)$ sublinear convergence has been established for softmax PG with constant step size in [18]. The analysis therein is essentially based on the smoothness and the gradient domination property of the value function, and thus requires the step size to be sufficiently small. In addition, an adaptive step size based on the norm of the gradient has been proposed in [16], with which softmax PG can achieve linear convergence. In this section, we will establish the $O(1/k)$ sublinear convergence of softmax PG to any constant step size. Additionally, a new adaptive step size selection rule will also be introduced.
In contrast to the optimization analysis framework adopted in [18], the idea of analysis here is more elementary and similar to the one used in Section 3. That being said, the extension is by no means trivial since the update of softmax PG is substantially different with that of PPG. In particular, we will establish the improvement lower bound of $\mathcal{T}^{k+1}V^{k}(s)-V^{k}(s)$ for softmax PG in terms of a function of $\max_{a}|\hat{A}^{k}_{s,a}|$ in contrast to $\max_{a}A^{k}_{s,a}$ due to the existence of $\pi_{s,a}^{k}$ in the exponent in (4.2). Moreover, as noted after Lemma 3.1, the establishment of the improvement lower bound in terms of $\max_{a}A^{k}_{s,a}$ for PPG relies heavily on the explicit formula for the projection which is not applicable for softmax PG. Instead, a complete different route is adopted to establish the improvement lower bound for softmax PG. Moreover, the improvement upper bound is also established which enables us to show that the $O(1/k)$ sublinear convergence rate is tight for softmax PG with any constant step size.
@@ 4.1 Improvement Lower and Upper Bounds
We begin with a novel identity that plays an important role in the establishment of the improvement lower and upper bounds.
@@
Noticing that $\max_{a}|\hat{A}^{k}_{s,a}|\leq\frac{1}{1-\gamma}$, together with the fact
$\exp\left(\frac{2\eta_{k}}{1-\gamma}x\right)-1\leq\left(\exp\left(\frac{2\eta_{k}}{(1-\gamma)^{2}}\right)-1\right)(1-\gamma)\,x,\quad\mbox{for }x\in\left[0,\frac{1}{1-\gamma}\right],$
one further has
$\sum_{a}\pi^{k+1}_{s,a}A^{k}_{s,a}\leq\left(\exp\left(\frac{2\,\eta_{k}}{(1-\gamma)^{2}}\right)-1\right)(1-\gamma)|\mathcal{A}|\left(\max_{a}|\hat{A}^{k}_{s,a}|\right)^{2}.$
@@ 4.2 Convergence Results
We first present the global convergence of softmax PG for any step size sequence $\eta_{k}$ as long as $\inf_{k}\eta_{k}\geq\alpha$ for some $\alpha>0$. The proof overall follows the route in [1] and is presented in Appendix B where we only point out how to generalize the proof in [1] to any step size sequence.
@@
In particular, if the optimal policy is unique, i.e., $|\mathcal{A}_{s}^{*}|=1$, one has
@@
$V^{*}(\rho)-V^{k}(\rho)\leq\frac{1}{k}\frac{1}{(1-\gamma)^{3}}\frac{2|\mathcal{A}|}{\kappa^{2}}\left\|\frac{d_{\rho}^{*}}{\rho}\right\|_{\infty}\left\|\frac{1}{\mu}\right\|_{\infty},$
which has better dependency on $1-\gamma$.
@@
The proof is completed by using Theorem 2.11 and noting that $\left\|\frac{d^{\xi^{k}}_{\rho}}{\rho}\right\|^{-1}\geq\tilde{\rho}$. ∎
@@
It follows easily from (4.5) and the performance difference lemma that $\sum_{a\in\mathcal{A}}{\pi_{s,a}^{k+1}A_{s,a}^{k}}\geq 0$ and $\underset{k\rightarrow+\infty}{\lim}|\hat{A}_{s,a}^{k}|=0$. Thus, $\eta_{k}$ is lower bounded. The application of Theorem 4.4 also implies the global convergence of softmax PG with the adaptive step size and thus $\kappa>0$. It is worth noting that the adaptive step size $\eta_{k}={\eta}/{\left\|\nabla_{\theta}V^{\pi_{\theta_{k}}}\left(\mu\right)\right\|_{2}}$ is considered in [17] and the following linear convergence rate has been established for softmax PG,